(25). 3. Find ilt) 61 422 21 to ses Zov ately T**
Answer #1

t=0 62 한 42 Mo 20v I + Ve vi F :25 before switching, there is no Factive supply.so Vcco=0 After switching the Sa domain circu2 val 35+2] 90 65 va = 20 s (35+2) Vat 2013 s(s+212) valta vid vaess? Vais) 2013 s(s+213) B S S+ 2/3 Alo = 10 B1-3 -10 3 Va(s20 -O Ils) = 35 s+ 3 ut) - 20g utt) amps 10023+ i sta CS Scanned with CamScanner